1.     Drive the AD (Aggregate Demand) curve using the following: IS curve is given as Y = 20XX-100i, LM1 is Y= 1000+25i (when P = 1) and LM2 is Y = 500+25i (when P =2), where XX is the last two digits of your student ID number. Show the derivation in (interest rate-income) and (price level-income) spaces. (You may insert a snapshot of the graphs if drawn manually).


1
Expert's answer
2021-05-05T13:33:44-0400

"solution"

The student ID number being S1117895. Therefore the last two digits are 95.

Given IS CURVE is :

"Y=20XX-100i\\
substitute\ for \ XX \ above:\\
therefore\ the \ IS \ curve \ is \ :\\
 Y=2095-100i"


Given :"LM_1 \ is \ Y=1000+25i\ (P=1)\\"

and "LM_2 \ is \ Y=500+25i\ (P=2)"


"IS : Y=2095-100i\\
LM_1:Y=1000+25i\\
IS=LM\\
2095-100i=100+25i\\
2095-1000=25i+100i\\
1095=125i\\
i=\frac{219}{125}=8.76\\
i=8.76"


"IS:Y=2095-100i\\
LM_2:Y=500+25i\\
IS=LM\\
2095-100i=500+25i\\
2095-500=25i+100i\\
1595=125i\\
i=\frac{319}{25}=12.76\\
i=12.76"




"Y^*_1=2095-(100\times 8.76)=1219\\
Y^*_2=2095-(100\times12.76)=819"


Derivation ="1219-819=\$400"


price level income :"Y_1=1000+(25\times1)=1025\\"

"Y_2=500+(25\times2)=550"


Derivation="1025-550=\$475"